OpenStudy (anonymous):

(5/7)power3/13/(5/7)power -10/13 five by seven whole pe power three by thirteen divided by five by seven whole pe power minus ten by thirteen siplify it

OpenStudy (anonymous):

Both the powers will be added.. \[3/13 +10/13 = 13/13 = 1..\] So, the answer will become 5/7..